Samples of size 25 are selected from a population with a mean of 40 and a standard deviation of 7.5 what is the mean of error of the sampling distribution of sample means

1
Expert's answer
2022-05-25T08:49:23-0400

Given: "n=25, \mu=40, \sigma=7.5"

"E(\=x)=\mu=40" , because the mean of the sampling distribution of the sample mean is equal to the population mean.

Standard deviation of the sampling distribution:

"\sigma_{\=x}=\frac{\sigma}{\sqrt{n}}=\frac{7.5}{5}=1.5"